# Disentangling CLIP for Multi-Object Perception

### v3

@@ Abstract
Vision-language models like CLIP excel at recognizing the single, prominent object in a scene. However, they struggle in complex scenes containing multiple objects. We identify a fundamental reason behind this limitation: VLMs features space exhibits significant semantic entanglement, where features of one class contain substantial information about other unrelated classes, a phenomenon we term mutual feature information (MFI). This entanglement becomes evident during class-specific queries, as unrelated objects are activated alongside the queried class. To address this limitation, we propose DCLIP, a framework that disentangles CLIP features using two complementary objectives: a novel MFI Loss that orthogonalizes the text (class) features to reduce inter-class similarity, and the Asymmetric Loss (ASL) that aligns image features with the disentangled text features. Our experiment demonstrates that DCLIP reduces inter-class feature similarity by 30% compared to CLIP, leading to significant performance gains on multi-label recognition (MLR) and zero-shot semantic segmentation (ZS3). In MLR, DCLIP outperforms SOTA approaches on VOC2007 and COCO-14 while using 75% fewer parameters, and surpasses SOTA ZS3 methods by 3.4 mIoU on VOC2012 and 2.8 mIoU on COCO-17. These results establish feature disentanglement as a critical factor for effective multi-object perception in vision-language models.
@@ 1 Introduction
Vision-language models (VLMs) such as CLIP have emerged as powerful tools for understanding visual content through natural language supervision. CLIP trains on a massive dataset of 400 million image-text pairs and demonstrates impressive performance in recognizing the salient object in the image, retrieving similar images from large datasets, and answering image-related natural language queries. However, an important question arises: do these impressive capabilities transfer when CLIP processes images containing multiple objects? As illustrated in Fig. 1, CLIP often struggles in such scenarios, failing to accurately recognize and localize all the objects present in the image. In this work, we investigate and identify the causes of this limitation and propose a framework that enables VLMs to handle complex images with multiple objects.
Figure 1: Illustration of Feature Entanglement in CLIP. Given the query “Human”, CLIP exhibits feature entanglement, spuriously activating regions corresponding to other objects (dogs and horse). In contrast, DCLIP (ours) accurately focuses only on the human. The rightmost plot quantifies this improvement through reduced cosine similarity between class features. Takeaway: By significantly decreasing inter-class similarity (Human-Horse: 0.84 $\rightarrow$ 0.42 [50% $\downarrow$], Human-Dog: 0.80 $\rightarrow$ 0.28 [65% $\downarrow$], DCLIP yields more disentangled features, crucial for precise multi-object perception.
@@
In our investigation, we systematically analyze CLIP’s features and identify two key factors contributing to this limitation. First, the spatial pooling operation in the visual encoder’s final layer, while sufficient for identifying the prominent object, eliminates crucial spatial information needed to recognize and locate multiple distinct objects in an image. Second, and more importantly, we discover significant entanglement between class features in the joint vision-language space, a phenomenon we term mutual feature information (MFI). This entanglement becomes apparent during class-specific queries: regions containing unrelated objects activate alongside the target class. For eg, as illustrated in Fig. 1, “dog” and “horse” regions also activate when we query “human.” This activation pattern strongly correlates with the high similarity scores between class features (0.84 for human-horse and 0.80 for human-dog). We extend this analysis to the classes in VOC [11] and COCO [19] datasets, where we observe average feature similarities of 0.77 in VOC and 0.69 in COCO (Tab. 5), confirming substantial feature entanglement in CLIP’s feature space.
@@
To address the feature entanglement in VLMs, we propose DCLIP, a framework that disentangles class features using two complementary objectives. We draw inspiration from the redundancy reduction principle [2], which states that sensory systems (eg. brain) recode input information such that redundancy is minimized without losing useful information, and extend it to the vision-language domain to reduce MFI. While previous approaches have focused on architectural modifications [30, 18, 3] or prompt engineering [27, 23] to adapt VLMs for multi-object settings, they do not address the fundamental feature entanglement problem. DCLIP targets this root cause through our novel MFI Loss that explicitly disentangles text features by orthogonalizing them to reduce inter-class similarity, while the Asymmetric Loss (ASL) ensures proper cross-modal alignment by projecting image features to align with the disentangled text features. This joint training using MFI loss (separates text features) and ASL loss (aligns text and image features) results in a semantically disentangled feature space that significantly improves multi-object perception capabilities.
@@
We evaluate DCLIP on multi-label recognition (MLR) and zero-shot semantic segmentation (ZS3). using established benchmarks: COCO-14 and VOC2007 [11] for MLR, and VOC2012 [11] and COCO-17 [19] for ZS3. Our experimental results demonstrate that DCLIP reduces inter-class feature similarity by an average of 30% compared to CLIP across these datasets, leading to favorable MLR performance over SOTA methods on VOC2007 and the challenging COCO-14 dataset, while requiring 75% fewer parameters. For ZS3, DCLIP surpasses SOTA VLM methods, demonstrating that reducing mutual feature information (MFI) is crucial for multi-object perception.
@@
  • We identify a critical limitation of VLMs for multi-object perception: mutual information between class features (MFI) negatively impacts recognition and localization performance
  • We propose DCLIP, a framework that disentangles class features through our novel MFI and the ASL loss, creating a disentangled feature space while preserving image-text alignment
  • The disentangled feature space is suitable for both MLR and ZS3, and outperforms SOTA on both tasks while requiring 75% fewer training parameters (trained only for MLR)
@@ 2 Related Work
Vision-Language Models for Multi-Object Perception. Vision-language models (VLMs) trained with contrastive losses [22, 16] are challenging to adapt for multi-object settings for two reasons: (1) their reliance on global feature aggregation, which ignores local information. (2) Using the softmax operation in their training loss biases them toward single-object settings.
@@
Recognition. Early efforts to adapt VLMs for recognition centered on learning prompts as classifiers for visual features [31]. These methods were extended to multi-label settings by learning multiple prompts for each class [27, 15, 23]. Subsequent works incorporated co-occurrence information to make predictions interdependent [9, 24]. In contrast, our approach does not rely on prompt learning or co-occurrence modeling during pre-training. Furthermore, our features are adaptable to tasks beyond multi-label recognition.
@@
Localization. Early approaches addressed localization by training image segmentation models and using VLMs to label the segmented regions [17]. Later methods introduced pre-training setups that combined vision-language alignment with mask distillation to enhance localization [10]. Recent works adapted features for localization without additional training by leveraging the spatial properties preserved in the value projection of CLIP’s transformer-style aggregation [30]. CLIP Surgery [18] identified consistent noisy activations across classes and reduced them by subtracting average features from class-specific features [18], though the cause of these activations remains unclear. GEM generalized this concept to vision transformers [3]. We use the finding that value projection preserves spatial information. We further improve value projection by disentangling class features.
@@
Recoding information. Shannon proposed that optimal information transmission involves designing codes with minimum entropy [26]. The redundancy reduction principle extended this idea to neuroscience, suggesting that sensory systems recode information to reduce redundancy with minimal loss [2]. This principle has since been applied to many recent works, including image compression [1] and more popularly in representation learning [21, 5, 29, 14, 13, 6]. While our loss function shares structural similarities with representation learning methods (a similarity and contrastive term), our method differs as follows: (1) DCLIP uniquely refines pre-trained VLM features, directly manipulating an existing, semantically rich space rather than learning representations from scratch (2) Crucially, our MFI loss operates on the fixed set of class text embeddings to directly reduce inter-class semantic similarity. This contrasts fundamentally with instance-discriminative contrastive losses that rely on intra-sample invariance (3) DCLIP is adapted for multi-object perception in VLMs, a setting prior works do not address.
@@
Figure 3: DCLIP Overview. Given image and class names in the dataset, CLIP extracts image and text features, which are projected by respective projectors to a disentangled space while preserving local image information. To reduce mutual feature information (MFI) between class features, we propose MFI loss that enforces the self-similarity matrix of projected text features to approximate an identity matrix, effectively reducing inter-class feature dependencies (Sec. 3.3). To align the image and the separated text features we the ASL loss in a multi-label recognition setup (Sec. 3.4). This setup aggregates the projected image and text features to obtain predicted logits (Sec. 3.4). The predicted logits are trained with ground truth labels using the widely used asymmetric loss (ASL) [25]. Our training loss combines the ASL and MFI loss, the only trainable components are the projectors. During inference, we freeze CLIP’s image and text encoders, along with our projectors for multi-label recognition and zero-shot semantic segmentation tasks.
@@ 3.1 Analysis of Feature Entanglement in CLIP
Before we introduce DCLIP, we first analyze the feature entanglement phenomenon in CLIP that limits its multi-object perception capabilities. Our examination of pairwise cosine similarity between class text features reveals unexpectedly high inter-class similarities in CLIP’s feature space, with average values of 0.77 for VOC and 0.69 for COCO classes (Fig. 2). Strikingly, even semantically distinct categories like “human” and “horse” show similarity scores as high as 0.84, far exceeding what their semantic relationship would suggest. This feature entanglement manifests visually when querying for specific classes in complex scenes. As shown in Fig. 1, when querying for “human,” CLIP erroneously activates “dog” and “horse” regions, a direct consequence of their entangled feature representations. The distribution of cosine similarity in Fig. 4 confirms this is not an isolated issue but rather a pervasive problem affecting the majority of class pairs. We attribute this entanglement to CLIP’s contrastive training objective, which optimizes for global image-text alignment without explicitly constraining separation between different class features. Additionally, the global pooling operation mixes features across the entire image, which further exacerbates feature entanglement. Our analysis establishes a clear correlation between feature entanglement reduction and improved performance on multi-object perception (Fig. 6). As mutual feature information (MFI) decreases, both multi-label recognition and semantic segmentation performance consistently improve, confirming that addressing feature entanglement is crucial for enhancing CLIP’s multi-object perception. This insight motivates our DCLIP framework, which maintains CLIP’s rich semantic knowledge while explicitly reducing mutual information between class features to improve multi-object perception.
@@ 3.3 MFI Loss
We design our projected feature space to reduce mutual feature information (MFI) between class features. To reduce MFI, we require isolated inputs for each class, as MFI measures the information shared between features of these isolated inputs. For images with multiple objects, obtaining isolated class inputs is challenging. For eg, an image containing both a horse and a dog would require computationally intensive segmentation methods to separate these objects. In contrast, text inputs are inherently isolated in our framework, as we already have individual class names (e.g., “horse” and “dog”) available in the dataset. We leverage this natural isolation of text inputs by extracting text features, projecting them into our shared space to get $\mathbf{t^{\prime}}$, and applying our proposed MFI loss to reduce information overlap between classes. Instead of using computationally expensive object segmentation for images, we propose an elegant solution that aligns these separated text features with image features using ASL loss (Sec. 3.4). Our MFI loss function is formulated as:
@@
where $\mathbf{S}$ is the self-similarity matrix and is given by $\mathbf{S}_{ij}=\mathbf{t^{\prime}}_{i}^{\top}\mathbf{t^{\prime}}_{j}$ where $\mathbf{t^{\prime}}_{i},\mathbf{t^{\prime}}_{j}$ are the $i$-th and $j$-th column vectors of $\mathbf{t^{\prime}}$ (i.e., $\mathbf{t^{\prime}}_{i},\mathbf{t^{\prime}}_{j}\in\mathbb{R}^{1\times d^{\prime}}$). In this formulation, $\lambda$ is the hyperparameter that addresses the imbalance in the loss arising from the larger number of MFI reduction terms in $\mathbf{S}$ compared to the collapse prevention terms. The MFI loss minimizes the inter-class similarity $\mathbf{S}_{ij}$ ($i\neq j$) while simultaneously preserving high intra-class $\mathbf{S}_{ii}$ to prevent feature collapse. We provide the detailed connection of MFI to the Information Bottleneck principle in Appendix Sec. A.
(a) VOC Dataset
(b) COCO Dataset
@@ 3.4 Image-Text Alignment with ASL
MLR Formulation. MLR task involves identifying the subset of classes $\mathcal{C}_{i}\subseteq\{C_{1},C_{2},\ldots,C_{N}\}$ associated with the image $\mathbf{x}_{i}$. The goal is to learn a mapping function $g:\mathbf{x}_{i}\rightarrow\{-1,1\}^{N}$, that maps input images to $1$ if the class is present and $-1$ if the class is absent in the image.
We align $\mathbf{t^{\prime}_{i}}$ and $\mathbf{z^{\prime}_{i}}$ using the ASL loss in the MLR setup, eliminating the need for image segmentation. For each location $(h,w)$ in $(\mathbf{z^{\prime}}_{i})$, we detect the presence or absence of a class $j$, by computing the cosine similarity with positive text features ($\mathbf{t^{\prime}_{j,+}}$) and negative text features ($\mathbf{t^{\prime}_{j,-}}$). Higher similarity with the positive text features indicates the presence of the class, while a higher similarity with the negative text features indicates its absence. This process generates class-specific positive and negative similarity maps across the image. Following [27, 24, 23], we apply softmax to these maps to increase focus on local regions containing the class. See Appendix B for details. We then aggregate across the local regions to obtain the logits ${p_{i}}$ for the image. We train the setup with the widely used Asymmetric Loss function (ASL) [25], which addresses the significant imbalance between negative and positive examples in multi-label dataset. The ASL loss is given by:
@@
where $p_{i}^{j}$ represents the corresponding prediction associated with label $y_{i}^{j}$, $i$ represents the image and $j$ represents the class. $p_{i,\delta}^{j}=\max(\hat{y}-\delta,0)$, with $\delta$ is the shifting parameter defined in ASL.
Table 1: Comparison on multi-label recognition (MLR). We compare the performance (mAP) and training efficiency (parameters, GPU hours) of our approach with SOTA VLM-based MLR methods on VOC2007 and COCO-14 datasets. DCLIP performs favorably over SOTA on VOC2007, and on the challenging COCO dataset, it outperforms SOTA while requiring only one-fourth of the parameters. The reported mAP values represent the mean and standard deviation over 5 runs of DCLIP. red and blue indicate the best and the second best performance.
Methods VOC2007 COCO-14
Param($\downarrow$) GPU hrs($\downarrow$) mAP($\uparrow$) Param($\downarrow$) GPU hrs($\downarrow$) mAP($\uparrow$)
DualCoOp [27] 0.3M 3.6 94.2 1.3M 16 83.6
SCPNet [9] - $>$3.6 94.3 3.4M 26 84.4
TAI-DPT [12] $>$0.3M - - $>$1.3M - 84.5
DualCoOp++ [15] 0.4M - 94.9 1.5M - 85.1
MLR-GCN [24] 0.4M 3.6 94.4 1.3M - -
PositiveCoOp [23] 0.2M 3 94.4 0.8M 15 84.7
Ours 0.4M 3 95.2 $\pm$ 0.2 0.4M 13 85.4 $\pm$ 0.2
Training. Our training objective is composed of two components: (1) mutual feature information loss that enforces the separation between class text features and (2) Asymmetric loss function [25], designed for MLR that aligns the image features and text features to obtain predictions for an image.
@@
where $\alpha$ controls the relative importance of the two objectives.

### v4

# Efficiently Disentangling CLIP for Multi-Object Perception

@@ Abstract
Vision-language models like CLIP excel at recognizing the single, prominent object in a scene. However, they struggle in complex scenes containing multiple objects. We identify a fundamental reason for this limitation: VLM feature space exhibits excessive mutual feature information (MFI), where the features of one class contain substantial information about other, unrelated classes. This high MFI becomes evident during class-specific queries, as unrelated objects are activated alongside the queried class. To address this limitation, we propose DCLIP, an efficient framework that learns an optimal level of mutual information while adding only minimal learnable parameters to a frozen VLM. DCLIP uses two complementary losses: a novel MFI Loss that regulates class feature similarity to prevent excessive overlap while preserving necessary shared information, and the Asymmetric Loss (ASL) that aligns image features with the disentangled text features. Through this disentanglement, DCLIP reduces excessive inter-class similarity by 30%. On multi-label recognition, DCLIP performs favorably over SOTA approaches on VOC2007 and COCO-14 while using 75% fewer training parameters. For zero-shot semantic segmentation, it shows improved performance across six benchmark datasets. These results highlight the importance of feature disentanglement for multi-object perception in VLMs.
@@ 1 Introduction
Vision-language models (VLMs) such as CLIP have emerged as powerful tools for understanding visual content through natural language supervision. CLIP trains on a massive dataset of 400 million image-text pairs and demonstrates impressive performance in recognizing the salient object in the image, retrieving similar images from large datasets, and answering image-related natural language queries. However, an important question arises: do these impressive capabilities transfer when CLIP processes images containing multiple objects? As illustrated in Fig. 1, CLIP often struggles in such scenarios, failing to accurately recognize and localize all the objects present in the image. In this work, we investigate and identify the causes of this limitation and propose a framework that enables VLMs to handle complex images with multiple objects efficiently.
@@
In our investigation, we analyze CLIP’s features and identify two key factors contributing to this limitation. First, the spatial pooling operation in the visual encoder’s final layer, while sufficient for identifying the prominent object, eliminates crucial spatial information needed to recognize and locate multiple distinct objects in an image. Second, and more importantly, we discover excessive entanglement between class features in the vision-language space, which we term mutual feature information (MFI). While some degree of feature similarity is useful for capturing broad semantic relationships (e.g., "dog" and "horse" share features as four legged animals), we find CLIP’s features are excessively entangled. This entanglement (high MFI) becomes apparent during class-specific queries as illustrated in Fig. 1, where “dog” and “horse” regions also activate when we query “human.” This activation pattern strongly correlates with the high similarity scores between class features (0.84 for human-horse, 0.80 for human-dog). We extend this analysis to the classes in VOC (Everingham et al., 2010) and COCO (Lin et al., 2014), where we observe average feature similarities of 0.77 in VOC and 0.69 in COCO (Tab. 4.2), confirming excessive feature entanglement in CLIP’s feature space.
Figure 1: Illustration of Feature Entanglement in CLIP. Given the query “Human”, CLIP exhibits high feature entanglement, spuriously activating regions corresponding to other objects (dogs and horse). In contrast, DCLIP (ours) accurately focuses only on the human. The rightmost plot quantifies this improvement through reduced cosine similarity between class features. Takeaway: By regulating inter-class similarity (Human-Horse: 0.84 $\rightarrow$ 0.42 [50% $\downarrow$], Human-Dog: 0.80 $\rightarrow$ 0.28 [65% $\downarrow$], DCLIP yields more disentangled features, crucial for precise multi-object perception.
@@
To address the feature entanglement in VLMs, we propose DCLIP, a lightweight framework that disentangles class features using two complementary objectives. We draw inspiration from the redundancy reduction principle (Barlow et al., 1961), which states that sensory systems (eg. brain) recode input information such that redundancy is minimized without losing useful information, and extend it to the vision-language domain to reduce MFI. While previous approaches have focused on architectural modifications (Zhou et al., 2022a; Li et al., 2025; Bousselham et al., 2024) or prompt engineering (Sun et al., 2022; Rawlekar et al., 2025) to adapt VLMs for multi-object settings, they do not address the fundamental feature entanglement problem and often come with significant computational overhead. In contrast, DCLIP explicitly targets this root cause while adding only minimal learnable parameters on top of frozen VLMs, making it both effective and efficient. Our novel MFI Loss orthogonalizes text features to regulate inter-class similarity, preventing excessive overlap while preserving necessary shared information. Meanwhile the Asymmetric Loss (ASL) ensures proper cross-modal alignment. This joint training (MFI + ASL) produces an optimally disentangled feature space that significantly improves multi-object perception capabilities at low cost.
@@
We evaluate DCLIP on multi-label recognition (MLR) and zero-shot semantic segmentation (ZS3). using established benchmarks: COCO-14 and VOC2007 (Everingham et al., 2010) for MLR. Importantly, for ZS3 evaluation, we use DCLIP projectors from MLR (trained on COCO-14) with image-level labels and evaluate it on six diverse datasets without any local annotations or fine-tuning. Our experimental results demonstrate that DCLIP reduces inter-class feature similarity by an average of 30% compared to CLIP across these datasets, leading to favorable MLR performance over SOTA methods on VOC2007 and the challenging COCO-14 dataset, while requiring 75% fewer parameters. For ZS3, DCLIP surpasses SOTA VLM methods, showing that reducing mutual feature information (MFI) is crucial for multi-object perception and can be achieved efficiently.
@@
  • We identify that excessive mutual information between class features (MFI) is the bottleneck in VLMs’ multi-object perception, leading to spurious cross-class activations
  • We propose DCLIP, an efficient framework that regulates mutual information between classes through novel MFI (grounded in information bottleneck principle) and ASL losses, creating disentangled features while preserving image-text alignment
  • We demonstrate DCLIP’s feature improvements through two tasks: trained only for multi-label recognition with 75% fewer trainable parameters, it achieves strong MLR performance and outperforms prior work on six zero-shot semantic segmentation benchmarks
@@ 2 Related Work
Vision-Language Models for Multi-Object Perception. Vision-language models (VLMs) trained with contrastive losses (Radford et al., 2021) are challenging to adapt for multi-object settings for two reasons: (1) Their reliance on global feature aggregation, which ignores local information. (2) The softmax operation in their training loss biases them toward single-object settings.
@@
Recognition. Early efforts to adapt VLMs for recognition centered on learning prompts as classifiers for visual features (Zhou et al., 2022b). These methods were extended to multi-label settings by learning multiple prompts for each class (Sun et al., 2022; Hu et al., 2023). Subsequent works incorporated co-occurrence information to make predictions interdependent (Rawlekar et al., 2024). In contrast, our approach does not rely on prompt learning or co-occurrence modeling during pre-training. Furthermore, our features are adaptable to tasks beyond MLR.
@@
Localization. Early approaches addressed localization by training image segmentation models and using VLMs to label the segmented regions (Kirillov et al., 2023). Later methods introduced pre-training setups that combined vision-language alignment with mask distillation to enhance localization (Dong et al., 2023). Recent works adapted features for localization without additional training by leveraging the spatial properties preserved in the value projection of CLIP’s transformer-style aggregation (Zhou et al., 2022a). CLIP Surgery (Li et al., 2025) identified consistent noisy activations across classes and reduced them by subtracting average features from class-specific features (Li et al., 2025), though the cause of these activations remains unclear. In contrast, we provide a principled analysis identifying high mutual feature information (MFI) as the cause of poor multi-object perception, and propose a theoretically grounded solution through our MFI loss.
@@
Recoding information. Shannon proposed that optimal information transmission involves designing codes with minimum entropy (Shannon, 1948). The redundancy reduction principle extended this idea to neuroscience, suggesting that sensory systems recode information to reduce redundancy with minimal loss (Barlow et al., 1961). This principle has since been applied to many recent works, including image compression (Ballé et al., 2016) and more popularly in representation learning (Oord et al., 2018; Chen et al., 2020; Zbontar et al., 2021; Henaff, 2020; He et al., 2020; Chen & He, 2021). While our loss function shares structural similarities with representation learning methods (a similarity and contrastive term), our method differs as follows: (1) DCLIP uniquely refines pre-trained VLM features, directly manipulating an existing, semantically rich space rather than learning representations from scratch (2) Crucially, our MFI loss operates on the fixed set of class text embeddings to directly reduce inter-class semantic similarity. This contrasts fundamentally with instance-discriminative contrastive losses that rely on intra-sample invariance (3) DCLIP is adapted for multi-object perception in VLMs, a setting prior works do not address.
@@
Figure 3: DCLIP Overview. Given image and class names in the dataset, CLIP extracts image and text features, which are projected by respective projectors to a disentangled space while preserving local image information. To reduce mutual feature information (MFI) between class features, we propose MFI loss that enforces the self-similarity matrix of projected text features to approximate an identity matrix, effectively regulating inter-class feature dependencies (Sec. 3.3). To align the image and the separated text features, we use the ASL loss in a multi-label recognition setup (Sec. 3.4). This setup aggregates the projected image and text features to obtain predicted logits (Sec. 3.4). The predicted logits are trained with ground truth labels using the widely used asymmetric loss (ASL) (Ridnik et al., 2021). Our training loss combines the ASL and MFI loss, the only trainable components are the projectors. During inference, we freeze CLIP’s image and text encoders, along with our projectors for multi-label recognition and zero-shot semantic segmentation tasks.
@@ 3.1 Analysis of Feature Entanglement in CLIP
Before we introduce DCLIP, we first analyze the feature entanglement (high MFI) in CLIP that limits its multi-object perception capabilities. Our examination of pairwise cosine similarity between class text features reveals unexpectedly high inter-class similarities in CLIP’s feature space, with average values of 0.77 for VOC and 0.69 for COCO classes (Fig. 2). Strikingly, even semantically distinct categories like “human” and “horse” show similarity scores as high as 0.84, far exceeding what their semantic relationship would suggest. This feature entanglement manifests visually when querying for specific classes in complex scenes. As shown in Fig. 1, when querying for “human,” CLIP erroneously activates “dog” and “horse” regions, a direct consequence of their entangled feature representations. The distribution of cosine similarity in Fig. 4 confirms this is not an isolated issue but rather a pervasive problem affecting the majority of class pairs. We attribute this entanglement to CLIP’s contrastive training objective, which optimizes for global image-text alignment without explicitly constraining separation between different class features. Additionally, the global pooling operation mixes features across the entire image, which further exacerbates feature entanglement. Our analysis establishes a clear correlation between feature disentanglement and improved performance on multi-object perception (Fig. 7). As mutual feature information (MFI) decreases, both multi-label recognition and semantic segmentation performance consistently improve, confirming that regulating feature entanglement is crucial for enhancing CLIP’s multi-object perception. This insight motivates our DCLIP framework, which maintains CLIP’s rich semantic knowledge while explicitly reducing mutual information between class features to improve multi-object perception.
@@ 3.3 MFI Loss
We design our projected feature space to regulate mutual feature information (MFI) between class features. To regulate MFI, we require isolated inputs for each class, as MFI measures the information shared between features of these isolated inputs. For images with multiple objects, obtaining isolated class inputs is challenging. For eg, an image containing both a horse and a dog would require computationally intensive segmentation methods to separate these objects. In contrast, text inputs are inherently isolated in our framework, as we already have individual class names (e.g., “horse” and “dog”) available in the dataset. We leverage this natural isolation of text inputs by extracting text features, projecting them into our shared space to get $\mathbf{t^{\prime}}$, and applying our proposed MFI loss to reduce information overlap between classes. Instead of using computationally expensive object segmentation for images, we propose an elegant solution that aligns these separated text features with image features using ASL loss (Sec. 3.4). Theoretical grounding: Our MFI loss is derived from the Information Bottleneck (IB) principle. We extend the standard IB formulation to explicitly account for inter-class information sharing:
$\mathcal{IB}=I(Z_{i};X_{i})-\beta\left[I(Z_{i},Y_{i})-I(Z_{i},Y_{j\neq i})\right]$ (1)
where $I$ is the mutual information, $Z_{i}$, $X_{i}$, and $Y_{i}$ are the features, input, and target for class $i$. Under gaussian assumptions for high-dimensional entropy estimation (detailed proof in Appendix A), this reduces to minimizing cross-class correlations while preserving within-class structure. Our MFI loss:
@@
where $\mathbf{S}$ is the self-similarity matrix and is given by $\mathbf{S}_{ij}=\mathbf{t^{\prime}}_{i}^{\top}\mathbf{t^{\prime}}_{j}$ where $\mathbf{t^{\prime}}_{i},\mathbf{t^{\prime}}_{j}$ are the $i$-th and $j$-th column vectors of $\mathbf{t^{\prime}}$ (i.e., $\mathbf{t^{\prime}}_{i},\mathbf{t^{\prime}}_{j}\in\mathbb{R}^{1\times d^{\prime}}$). In this formulation, $\lambda$ is the hyperparameter that addresses the imbalance in the loss arising from the larger number of MFI reduction terms in $\mathbf{S}$ compared to the collapse prevention terms. The MFI loss minimizes the inter-class similarity $\mathbf{S}_{ij}$ ($i\neq j$) while simultaneously preserving high intra-class $\mathbf{S}_{ii}$ to prevent feature collapse.
@@ 3.4 Image-Text Alignment with ASL
MLR Formulation. MLR task involves identifying the subset of classes $\mathcal{C}_{i}\subseteq\{C_{1},C_{2},\ldots,C_{N}\}$ associated with the image $\mathbf{x}_{i}$. The goal is to learn a mapping function $g:\mathbf{x}_{i}\rightarrow\{-1,1\}^{N}$, that maps input images to $1$ if the class is present and 1-1 if the class is absent in the image.
We align $\mathbf{t^{\prime}_{i}}$ and $\mathbf{z^{\prime}_{i}}$ using the ASL loss in the MLR setup, eliminating the need for image segmentation. For each location $(h,w)$ in $(\mathbf{z^{\prime}}_{i})$, we detect the presence or absence of a class $j$, by computing the cosine similarity with positive text features ($\mathbf{t^{\prime}_{j,+}}$) and negative text features ($\mathbf{t^{\prime}_{j,-}}$):
$\mathbf{l}^{+}_{ij}[h,w]=\mathbf{z}_{i}[h,w]\cdot\mathbf{t}^{\prime}_{ij,+},\quad\mathbf{l}^{-}_{ij}[h,w]=\mathbf{z}_{i}[h,w]\cdot\mathbf{t}^{\prime}_{ij,-}$ (3)
Higher similarity with positive text features indicates class presence, while higher similarity with negative features indicates absence. Following (Sun et al., 2022; Rawlekar et al., 2024; 2025), we apply softmax to focus on relevant regions ($q$) and aggregate to obtain final logits ${p_{i}}=[\mathbf{p}^{+}_{i},\mathbf{p}^{-}_{i}]$:
$\mathbf{q}^{\pm}_{i}[h,w]=\frac{\exp(\mathbf{l}^{\pm}_{i}[h,w])}{\sum_{h^{\prime},w^{\prime}}\exp(\mathbf{l}^{\pm}_{i}[h^{\prime},w^{\prime}])},\quad\mathbf{p}^{\pm}_{i}=\sum_{h,w}\mathbf{q}^{\pm}_{i}[h,w]\cdot\mathbf{l}^{\pm}_{i}[h,w]$ (4)
See Appendix B for details. We train using Asymmetric Loss (ASL) (Ridnik et al., 2021). Here, $p_{i}^{j}$ is the prediction for label $y_{i}^{j}$, and $p_{i,\delta}^{j}=\max(\hat{y}-\delta,0)$ with shifting parameter $\delta$.
@@
Table 1: Comparison on multi-label recognition (MLR). We compare the performance (Precision and mAP) and training efficiency (parameters, GPU hours) of our approach with SOTA VLM-based MLR methods on VOC2007 and COCO-14 datasets. DCLIP performs favorably over SOTA on VOC2007, and on the challenging COCO dataset, it outperforms SOTA while requiring only one-fourth of the parameters. red and blue indicate the best and the second best performance.
Methods VOC2007 COCO-14 Param($\downarrow$) GPU hrs($\downarrow$) P($\uparrow$) mAP($\uparrow$) Param($\downarrow$) GPU hrs($\downarrow$) P($\uparrow$) mAP($\uparrow$) DualCoOp (Sun et al., 2022) 0.3M 3.6 80.5 94.2 1.3M 16 72.9 83.6 SCPNet (Ding et al., 2023) - >>3.6 94.3 3.4M 26 84.4 TAI-DPT (Guo et al., 2023) >>0.3M - - >>1.3M - 84.5 DualCoOp++ (Hu et al., 2023) 0.4M - 94.9 1.5M - 76.3 85.1 MLR-GCN (Rawlekar et al., 2024) 0.4M 3.6 81.1 94.4 1.3M - - PositiveCoOp (Rawlekar et al., 2025) 0.2M 3 81.7 94.4 0.8M 15 74.5 84.7 DCLIP (Ours) 0.4M 3 85.6 95.4 0.4M 13 80.1 85.6
Training. Our training objective is composed of two components: (1) mutual feature information loss that enforces the separation between class text features and (2) Asymmetric loss function (Ridnik et al., 2021), designed for MLR that aligns the image features and text features to obtain predictions for an image. Here, $\alpha$ controls the relative importance of the two objectives